# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients

### v2

@@ 4 Numerical Experiments
Verification of the Computational Complexity scaling: We verify the $\mathcal{O}(N)$ computational complexity claim next. We vary $N$ from $100$ to $30000$ and the time step is fixed at $\Delta t=10^{-3}$. Such range of mesh sizes is chosen to prevent the $O(r^{4})$ contributions of the ACS-GMRES solver from polluting the mesh-resolution scaling. Figure 5(a) displays $N$ vs. computational time for $10$ time steps of integration. We observe that the complexity remains linear with respect to $N$. The finite y-intercept corresponds to the $\mathcal{O}(r^{4})$ cost of the GMRES-ACS solve in 3D. Figure 5(b) displays the computational time required for the preconditioned GMRES-ACS solver vs. rank for a single time step, using BE with $N=10,000$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$. For this test, we allowed the basis to grow without truncation, recording the simulation time required for the GMRES solve at each basis augmentation iteration. The expected $\mathcal{O}(r^{4})$ scaling in 3D is found. These studies verify that the complexity scaling of our approach is $\mathcal{O}(Nr^{2}+r^{4})$ in 3D.
Figure 5: Example 4.1. (a) Demonstration of ${\cal O}(N)$ computational complexity of DIRK3 adaptive-rank integrator in 3D. (b) Demonstration of ${\cal O}(r^{4})$ computational complexity of the GMRES-ACS solver in 3D. The execution times were recorded using MATLAB’s timeit function.
@@
$f_{t}+\left[(1-x^{2})\cdot 2y\cdot-2z\cdot f\right]_{x}+\left[2x\cdot(1-y^{2})\cdot 2z\cdot f\right]_{y}+\left[-2x\cdot 2y\cdot 2(1-z^{2})\cdot f\right]_{z}=\nu\Delta f,\quad x,y,z\in[-1,1].$ (35)
We set the diffusion coefficient $\nu=1/500$. The initial condition is a bi-Gaussian function given by:
@@
This example demonstrates the algorithm’s ability to adapt to dynamic rank evolution. The algorithm performs nine xKrylov augmentations at the initial time step, two augmentations at the third and seventh time steps, and a single augmentation otherwise. Figure 6 shows rank evolution. The solution rank rises sharply in the $x$ and $y$ directions [Figures 6 (a)-(b)]. The solution rank in the $z$ direction evolves more gently [Figure 6 (c)] and is roughly one-third of that in the $x$ and $y$ directions. Figures 6 (d)-(e) showcase how the solution deforms according to the prescribed advection coefficients and diffuses slowly over time. Figure 7 shows snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$. As the flow-driven deformation increases the complexity of the solution structure, the rank grows to capture these features; when the diffusion process starts to dominate, the rank decreases as the solution smooths. This example shows that the low-rank assumption is not intrinsic to our proposed algorithm. Even when the rank becomes moderately large, the algorithm remains capable of capturing the underlying dynamics effectively. Naturally, the benefit of a low-rank approximation decreases as the solution approaches a full-rank structure.
Figure 6: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(c) Evolution of rank, xKrylov subspace before and after truncation, and number of xKrylov iterations over time for the $x$, $y$, and $z$ directions, respectively. (d)–(e) Solution profiles at $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$. The final simulation time is $T_{f}=10$.
Figure 7: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(f) snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$, corresponding to times $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$, respectively.
@@
The last example prescribes diffusion and advection coefficients that exactly balance each other, resulting in a non-trivial steady-state solution. We seek an equilibrium solution satisfying the steady-state condition:
@@
$\sigma_{1}^{1}(x)$ $=2x(1-3x+2x^{2}),\qquad\sigma_{1}^{2}(y)=y^{2}(1-y)^{2},\qquad\sigma_{2}^{1}(x)=x^{2}(1-x)^{2},\qquad\sigma_{2}^{2}(y)=2y(1-3y+2y^{2}).$
@@
Figure 8 shows the error of the numerical steady-state solution, reached using ten time steps, relative to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, for a fixed time step $\Delta t=1000$ ranges from $1,190$ to $36,076$ as the mesh is refined. The results are obtained by rescaling the numerical solution so that its normalization constant is the same as the analytical one. They confirm second-order spatial convergence and demonstrate the ability of our algorithm to take large timesteps to reach the correct steady state with expected asymptotic accuracy.
@@ 5 Conclusion
We have proposed a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients. Our algorithm is based on several key strategies: constructing dimension-wise basis from xKrylov subspaces, projecting the original matrix equation onto a reduced system to solve for a coefficient matrix, devising an effective Averaged-Coefficient Sylvester (ACS) preconditioner for the reduced system, and a computationally efficient procedure for the residual evaluation. This combined approach achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$, with $r$ being the rank of the solution during Krylov iteration process and $d$ the dimensionality, effectively mitigating the curse of dimensionality and making it comparable in efficiency to the constant-coefficient case [14]. Such complexity scaling is numerically verified with our extensive tests, with the maximal rank $r$ found to be independent of the 1D grid size $N$ for practical grid resolutions and convergence tolerances. Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
@@ Appendix A Discretization of the Diffusion Operator
  • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
@@ Appendix B Discretization of the Advection Operator
We discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction. To preserve linearity of our discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative. This discretization is monotonicity-preserving when the following condition is satisfied:
@@
where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity and $\phi_{\text{max}}$ is the maximum diffusion coefficient. This condition is practical in our formulation due to the very favorable scaling of our low-rank representation with mesh size and rank.
@@
  • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
@@ C.1 DIRK Scheme
$c_{s}$ $a_{s1}$ $a_{s2}$ldots $a_{ss}$
$b_{1}$ $b_{2}$ldots $b_{s}$
We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage. The Butcher tables for DIRK2 and DIRK3 are presented in Tables 3 and 3, respectively.
@@
$1$ $-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
$-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
@@
$\begin{split}\mathbf{F}^{(k)}&=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k}a_{k\ell}\mathbf{Y}_{\ell},\qquad k=1,2,\ldots,s,\end{split}$ (48a)
$\begin{split}\mathbf{Y}_{k}&=\mathscr{L}(\mathbf{F}^{(k)};t^{(k)}),\qquad t^{(k)}=t^{0}+c_{k}\Delta t,\qquad k=1,2,\ldots,s,\end{split}$ (48b)
@@
${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$ (49)
@@
$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s.$ (50)
@@ C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
[…]
$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$ (51)
@@
We follow [14] to perform high-order integration. The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${{\bf S}}_{1}$ through the stages. The proposed procedure is as follows:
@@
      $\tilde{Y}_{\ell}$ $=\frac{1}{a_{\ell\ell}\Delta t}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$
@@
      $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}_{1}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$ (55)

### v3

@@ 4 Numerical Experiments
Verification of the Computational Complexity and Memory Scalings: We verify the $\mathcal{O}(N)$ computational complexity and memory storage claims next. We vary $N$ from $100$ to $30000$ and the time step is fixed at $\Delta t=10^{-3}$. Figure 5(a) displays a profiling study of the algorithm, depicting the computational time of various components vs. $N$ obtained for $10$ time steps. We observe that only the xKrylov component scales with mesh refinement $N$, and does so linearly. Other components do not scale with $N$, consistently with the rank of the solution being bounded. Furthermore, the inner solve remains the dominant computational cost of the algorithm until, for grids of size $N\approx 20{,}000$, the xKrylov routine begins to dominate. This, in turn, highlights the importance of an efficient reduced-system solution strategy. Figure 5(b) displays the computational time required for the preconditioned GMRES-ACS solver vs. rank for a single time step, using BE with $N=10^{4}$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$. For this test, the xKrylov basis was allowed to continue to grow boundlessly, albeit with truncation applied at each augmentation step. The simulation time required for the GMRES solve was recorded at each basis augmentation. The expected $\mathcal{O}(r^{4})$ scaling in 3D is found. These studies verify that the complexity scaling of the approach is $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{4})$ in 3D. Figure 5(c) depicts the results of a memory scaling study demonstrating that the memory requirements also scale as $\mathcal{O}(N)$. The memory usage reported corresponds to the total size of all variables allocated within the solver function, as measured by MATLAB’s whos command at the end of the function call. A similar trend is expected (and observed in our simulations) in terms of memory storage: the inner tensor dominates the storage for coarser grids, while storing the xKrylov subspace dominates for finer grids.
Figure 5: Example 4.1. (a) Profiling study depicting the computational cost of various components of the algorithm, demonstrating overall ${\cal O}(N)$ computational complexity of DIRK3 adaptive-rank integrator in 3D. (b) Demonstration of ${\cal O}(r^{4})$ computational complexity of the GMRES-ACS solver in 3D. For this test only, we allow the rank to grow indefinitely by setting the outer residual tolerance, $\epsilon_{\texttt{tol}}$, to machine precision. (c) Demonstration of ${\cal O}(N)$ memory storage complexity of DIRK3 adaptive-rank integrator in 3D. The execution times were recorded using MATLAB’s timeit function. The memory usage was recorded using MATLAB’s whos command at the end of each time step, and the values reported correspond to the average memory usage across all time steps in megabytes.
To provide additional perspective on our results, we compare the performance of the adaptive-rank integrator for this example with a hypothetical optimal full-rank multigrid solver, which at best would scale as $\mathcal{O}(N^{d}\log N)$ [21]. Assuming that the low-rank and multigrid solvers are at parity for $N=100$ (Appendix E), Table 1 reports the theoretical wall-clock-time speedup of the adaptive-rank integrator vs. the full-rank multigrid solver. The table suggests vast efficiency gains from our approach vs. optimal multigrid for this problem, scaling as 1D with grid refinement and effectively mitigating the curse of dimensionality.
Table 1: Theoretical speedup of the adaptive-rank solver versus a hypothetical optimal multigrid solver for various 1D grid resolutions NN in a 3D domain.
1D grid size NN Wall-clock-time speedup in 3D
$1000$ $1446.1$
$4000$ $99229.3$
$10000$ $1418193.2$
@@
$f_{t}+\left[-4(1-x^{2})yzf\right]_{x}+\left[-4x(1-y^{2})zf\right]_{y}+\left[8xy(1-z^{2})f\right]_{z}=\nu\Delta f,\qquad x,y,z\in[-1,1].$ (35)
We set the diffusion coefficient $\nu=1/500$. The flow is solenoidal, and therefore a pure-vorticity flow. The initial condition is a bi-Gaussian function given by:
@@
This example demonstrates the algorithm’s ability to adapt to dynamic rank evolution. The algorithm performs nine xKrylov augmentations at the initial time step, two augmentations at the third and seventh time steps, and a single augmentation otherwise. Figure 6 shows rank evolution. The solution rank rises sharply in the $x$- and $y$-directions [Figures 6 (a)-(b)]. The solution rank in the $z$-direction evolves more gently [Figure 6 (c)] and is roughly one-third of that in the $x$ and $y$ directions. Figures 6 (d)-(e) showcase how the solution deforms according to the prescribed advection coefficients and diffuses slowly over time. Figure 7 shows snapshots of the solution in the $x\text{--}z$ plane at y=0.5y=-0.5. As the flow-driven deformation increases the complexity of the solution structure, the rank grows to capture these features; when the diffusion process starts to dominate, the rank decreases as the solution smooths. This example shows that the low-rank assumption is not intrinsic to the algorithm. Even when the rank becomes moderately large, the algorithm remains capable of capturing the underlying dynamics effectively. Naturally, the benefit of a low-rank approximation decreases as the solution approaches a full-rank structure.
From a computational perspective, the per-time-step cost breakdown is as follows: xKrylov construction and orthogonalization (10%), ACS-GMRES inner solve (47.5%), and residual evaluation (42.5%). The relatively small contribution of the cost of xKrylov construction is partly a consequence of the truncation mechanism, together with the effectiveness of the chosen projection spaces, which collectively keep the approximation spaces compact. On average, in the $x$-$y$ directions, the truncation reduces the subspace dimension by a factor of about $4.5$, from roughly $450$ to $100$. In the $z$ direction, the reduction amounts to a factor of approximately $3$, from about $150$ to $50$. On the other hand, the relatively large contribution of the residual evaluation to the simulation time originates in the poor scaling of the computational complexity with the number of GSE terms (eight in this example, including the RHS) and suggests that a more economical stopping criterion (as discussed in Section 3.4) should be used.
Figure 6: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(c) Evolution of rank, xKrylov subspace before and after truncation, and number of xKrylov iterations over time for the $x$, $y$, and $z$ directions, respectively. (d)–(e) Solution profiles at $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$. The red dots in (a)–(c) mark these time instants and correspond to the profiles shown in (d)–(e). The final simulation time is $T_{f}=10$.
Figure 7: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(f) snapshots of the solution in the $x\text{--}z$ plane at y=0.5y=-0.5, corresponding to times $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$, respectively.
@@
The last example prescribes diffusion and advection coefficients that exactly balance each other, resulting in a non-trivial steady-state solution. This test demonstrates the ability of the algorithm to take large timesteps stably to solve directly for the steady state with expected spatial asymptotic accuracy.
We seek an equilibrium solution satisfying the steady-state condition:
@@
$\sigma_{1}^{1}(x)$ $=2x(1-3x+2x^{2}),\hskip 18.49988pt\sigma_{1}^{2}(y)=y^{2}(1-y)^{2},\hskip 18.49988pt\sigma_{2}^{1}(x)=x^{2}(1-x)^{2},\hskip 18.49988pt\sigma_{2}^{2}(y)=2y(1-3y+2y^{2}).$
@@
Figure 8 shows the error of the numerical steady-state solution relative to the prescribed analytical steady state as a function of mesh size. The solution is obtained using ten time steps of size $\Delta t=1000$. The advection CFL number $\lambda_{A}$ ranges from $1,190$ to $36,076$ as the mesh is refined. The error is obtained by comparing the numerical and analytical solutions, with the numerical solution rescaled so that its normalization constant is the same as the analytical one. The results confirm second-order spatial convergence and the ability of the algorithm to solve directly for the correct steady state.
@@ 5 Conclusion
We propose a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients. We formulate the corresponding implicit system as a GSE, for which we follow a projection type approach, namely: constructing dimension-wise basis from xKrylov subspaces, projecting the original matrix equation onto a reduced system, and solving for the coefficient matrix. Our approach innovates in the way we construct the Krylov subspaces, and in the solution of the reduced system, which leverages GMRES effectively preconditioned with an effective Averaged-Coefficient Sylvester (ACS) approximation. The approach achieves a 1D-like computational complexity (of $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$, with $r$ being the rank of the solution during Krylov iteration process and $d$ the dimensionality) and memory storage [of $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$], effectively mitigating the curse of dimensionality when the solution rank is bounded and making it comparable in efficiency to the constant-coefficient case [14]. Such complexity scaling is numerically verified with extensive tests, with the maximal rank $r$ found to be independent of the 1D grid size $N$ for practical grid resolutions and convergence tolerances. Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
@@ Appendix A Discretization of the Diffusion Operator
  • $\mathbf{F}$: matrix of solution values, $\mathbf{F}]_{i,j}=f_{i,j}$.
@@ Appendix B Discretization of the Advection Operator
We discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction. To preserve linearity of the discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative. This discretization is monotonicity-preserving when the following condition is satisfied:
@@
where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity and $\phi_{\text{max}}$ is the maximum diffusion coefficient. This condition is practical in our formulation due to the very favorable scaling of the low-rank representation with mesh size and rank.
@@
  • $\mathbf{F}$: matrix of solution values, $\mathbf{F}]_{i,j}=f_{i,j}$.
@@ C.1 DIRK Scheme
$c_{s}$ $a_{s1}$ $a_{s2}$ $\ldots$ $a_{ss}$
$b_{1}$ $b_{2}$ $\ldots$ $b_{s}$
We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage. The Butcher tables for DIRK2 and DIRK3 are presented in Tables 4 and 4, respectively.
@@ 11 $1-\gamma$ $\gamma$
11 $\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
$\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
@@ 11 $\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
$\begin{split}\mathbf{F}^{(k)}&=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k}a_{k\ell}\mathbf{Y}_{\ell},\hskip 18.49988ptk=1,2,\ldots,s,\end{split}$ (48a)
$\begin{split}\mathbf{Y}_{k}&=\mathscr{L}(\mathbf{F}^{(k)};t^{(k)}),\hskip 18.49988ptt^{(k)}=t^{0}+c_{k}\Delta t,\hskip 18.49988ptk=1,2,\ldots,s,\end{split}$ (48b)
@@
${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\qquad\text{with}\qquad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\qquad k=1,2,\ldots,s.$ (49)
@@
$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\qquad k=1,2,\ldots,s.$ (50)
@@ C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
The high-order DIRK method features a stage-by-stage generalized-Sylvester-equation solve. The strategy follows closely that of the backward Euler adaptive-rank treatment, i.e., we construct an approximate $\mathscr$$\tilde${L}} used for both the basis construction of $U_{1}$ and $V_{1}$ and the preconditioning of the reduced system to solve $S_{1}$.
At each stage $k$ of the DIRK scheme, the corresponding DIRK formulation for the matrix differential equation from $t^{(0)}$ to $t^{(1)}$ requires solving an SE for the approximated operator $\mathscr$$\tilde${L}} by solving the $k$th-stage residual:
$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\qquad\text{with}\qquad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$ (51)
@@
We follow [14] to perform high-order integration. The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${{\bf S}}_{1}$ through the stages. The procedure is as follows:
@@
[…]
@@
[…]